The volume of 1.43 mol of gas at 63.0 kPa and 23.00 °C is ________ L.           
Answer #1

Answer:

given that,

Number of moles of gas (n) = 1.43 mol

Pressure (P) = 63.0 kpa = 0.621762 atm

Temperature (T) = 230c = 273 + 23 = 296 k

gas constant (R) = 0.0821 L-atm/deg/mol

Volume (V) = ?

Ideal gas equation, PV = nRT

V = (nRT) / P

= ( 1.43 \times 0.0821 \times 296 ) / 0.621762

= 55.89 L